OpenStudy (anonymous):

HELP PLSPLSPLSPLS INCLUDE STEPS. A fitness center has a membership fee of $125. Members only pay $5 per day to work out at the center. A nonmember pays $10 per day to work out. After how many work-out days is the total cost for members, including the membership fee, the same as the total cost for nonmembers

OpenStudy (anonymous):

member fee = 125 + 5d non member fee = 10d 125 + 5d = 10d 125 = 10d - 5d = 5d 25 = d 25 days i think

OpenStudy (anonymous):

Okay, lets set this up as an equation. Or two actually. You want to find the time when both members and non members pay the same amount. This can be done by setting up 125 + 5x = 10x. The variable (x) will be the same for both sides. X represents the amount of days. Now lets Get our like variables on the same side. So subtract 5x from 10x, which equals 5x. Your equation will now look like 125 = 5x. Divide 5 from each side to get X by its self, and you now have x = 25. Plug this back into you original equations to make sure. 125 + 5(25) = 250 10(25) = 250 So it would take 25 days for both members and non-members to pay the same amount.

OpenStudy (anonymous):

read tiffany's its much more detailed

hero (hero):

Tiffany go back to your question

OpenStudy (anonymous):

THANK YOU.